OpenStudy (anonymous):

whats the steps on coverting 9 3/4

OpenStudy (anonymous):

9X4=36 36+3=39 39/4

OpenStudy (anonymous):

indef. integral [ 9 + (3/4)] = 9x + (3/4)x Let x = 1. 9(1) + (3/4)(1) (9+(3/4))(1) Let A = 9 (A + (3/4))(1) A(4/4) = 4A/4 Substitute. ((4(9)/4)+(3/4))(1) ((36/4)+(3/4))(1) Let (36/4)+(3/4) = (36+3)/4 36+3 = 39 Substitute. (39 / 4)(1) = 39/4 Memorize these steps and you will be able to solve anything.